OpenStudy (anonymous):

(-1+5i)to the seventh power in rectangular form.

OpenStudy (anonymous):

write it as r[cost + isint] = re^(it) r = |z| and t = arctan(-5) z^7 = r^7 x e^(7it) = r^7 x (cos7t + isin7t]

OpenStudy (anonymous):

z = 26^(7/2) x [-7 + 35i]

OpenStudy (anonymous):

got it?